OpenStudy (anonymous):

plzz Equation M: y = x + 4 Equation P: y = 3x + 6 Which of the following is a step that can be used to find the solution to the set of equations? x = 3x + 6 x + 4 = 3x + 6 x + 6 = 3x + 4 x = 3x

OpenStudy (hayleymeyer):

we know that y = x + 4 and y = 3x + 6 so x+4=3x+6 so 3option is the answer :)